For each Molecule (1-11) Determine the A) hybridization, B) bond angle and C) whether the molecule is polar or nonpolar.

  1. BeF2
  2. NCl3
  3. SiS2
  4. CO(NH2)2
  5. SF 1-
  6. BCl3
  7. NO 1-
  8. SO 2-
  9. HCN
  10. H2CO
  11. ClF4+
Answer #1

AR Be fa valence é o be: - a L no. trobe auxf (A) Hybridisation : 2 -3 = o lone paus ano. of valence é 20 bonds by eacht . St(B) bond angle = 180° S = biss u=0 ☺ non - polar @ CNuala .. ☺ @ uto sp2 hybridisation of bonds = 3 no. of o (B) bond (C) angBulz- 8. S.No. 5 § - 3 valme é (voz) 0 VE 3 o bonds by 34 aroms . 3.= Stevic na = spa hybridisatia (6) bond angle = 120° B laNoo S. No. = 2+1 =3 lip o bond. ny bridisations = sp2 (11) Band angle = 120 (im) polar molecule Noe